OpenStudy (anonymous):

what is the common ratio of 6, 18, 54, 162

OpenStudy (anonymous):

3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

well, 6 *3 = 18 18* 3 = 54 54 * 3 = 162 what the person above me said